# DeepCRE: Revolutionizing Drug R&D with Cutting-Edge Computational Models

### 09 Mar 2024 v2

@@ Discussion
In this study, we have proposed the DeepCRE model and demonstrated its potential to revolutionize the drug R&D process. Through advancements in pretraining strategies, DeepCRE surpasses the existing SOTA models, achieving an average performance improvement of 17.7% in patient-level CRE, and a remarkable 5-fold increase in indication-level CRE. Moreover, DeepCRE has identified six drug candidates that show significantly greater effectiveness than a comparator set of two approved drugs in 5/8 CRC organoids tested. Besides, we have uncovered some pharmacodynamic and pharmacological insights into certain drug candidates, such as GSK1059615 and Dasatinib, which could inform further R&D process. Importantly, it is not merely about identifying one or two potential drugs, rather, it highlights the capability of DeepCRE to uncover a collection of drug candidates with enhanced therapeutic effects.
There are several future directions to explore. Firstly, leveraging Large Language Models (LLMs) for patient GEPs holds tremendous potential[41, 42]. Encoding GEPs using specialized genomic or multi-modal LLMs, which integrate gene language with clinical indicators and patient descriptions, could yield more comprehensive insights. Secondly, we advocate for expanding drug R&D beyond the traditional screening paradigm to include the generation paradigm. Generating models enables the active design of molecules with desired properties, offering benefits such as targeted design, cost and time savings, and fostering innovation by creating novel chemical entities. This approach reduces reliance on extensive screening of existing compounds and holds promise for developing breakthrough therapies with higher efficacy and lower failure rates during clinical trials. Finally, establishing a synergistic dry-wet-loop framework that integrates computational predictions with experimental validations can accelerate the drug R&D process. This integration facilitates seamless translation from pre-clinical research to clinical trials, ultimately benefiting patients in need.
@@ Methods
DeepCRE were employed to generate the DRP score matrix for the patients of COAD and READ, two subtypes of CRC. This matrix facilitated the ranking of drug candidates based on their efficacy across patient populations, as illustrated in Supplementary Figure 13a. Additionally, we conducted a comprehensive search for clinical records of these drug candidates within the ClinicalTrials database. For clarity and ease of comparison, the evaluation outcomes of DeepCRE alongside traditional methodologies were depicted in an XY plot (Fig. 5c). Notably, the analysis delineated the drug candidates into four distinct sets (A, B, C, and D). Sets A and B comprised drugs favored by DeepCRE predictions but with limited clinical trial evidence, whereas Sets C and D included drugs that were either approved or had substantial clinical trial involvement but were underestimated by DeepCRE predictions.

### 18 Mar 2024 v3

@@ Discussion
In this study, we have proposed the DeepCRE model and demonstrated its potential to transform the drug R&D process. Through advancements in pretraining strategies, DeepCRE surpasses the existing SOTA models, achieving an average performance improvement of 17.7% in patient-level CRE, and a remarkable 5-fold increase in indication-level CRE. Moreover, DeepCRE has identified six drug candidates that show significantly greater effectiveness than a comparator set of two approved drugs in 5/8 CRC organoids tested. Besides, we have elucidated some pharmacodynamic and pharmacological insights into certain drug candidates, such as GSK1059615 and Dasatinib, which could inform further R&D process. Importantly, it is not merely about identifying one or two potential drugs, rather, it highlights the capability of DeepCRE to systematically uncover a spectrum of drug candidates with enhanced therapeutic effects.
There are several future directions to explore. Firstly, leveraging Large Language Models (LLMs) for patient GEPs holds tremendous potential [41, 42]. By encoding GEPs with specialized genomic or multi-modal LLMs, which integrate gene languages with natural languages (clinical indicators and patient descriptions), we can derive more comprehensive insights. Such integration promises to revolutionize our understanding and approach to personalized medicine. Secondly, we propose a shift in drug R&D from the conventional screening methodology to a generative model-based approach. Generating models enables the active design of molecules with desired properties, offering benefits such as targeted design, cost and time savings, and fostering innovation by creating novel chemical entities[43]. By moving away from the exhaustive screening of existing compounds, this strategy presents a promising avenue for the discovery of groundbreaking therapies characterized by enhanced efficacy and reduced rates of failure in clinical trials. Finally, the establishment of a synergistic dry-wet-loop framework, integrating computational predictions with experimental validations, is critical for accelerating the drug R&D process. This harmonious integration promises a more efficient transition from pre-clinical research to clinical trials, directly translating to faster access to new treatments for patients. Such advancements underscore the importance of leveraging cutting-edge technologies and interdisciplinary approaches to overcome current limitations and fulfill the promise of next-generation medical breakthroughs.
@@ Methods
DeepCRE were employed to generate the DRP score matrix for the patients of COAD and READ, two subtypes of colorectal cancer (CRC). This matrix facilitated the ranking of drug candidates based on their efficacy across patient populations, as illustrated in Supplementary Figure 13a,b. Additionally, we conducted a comprehensive search for clinical records of these drug candidates within the ClinicalTrials database. For clarity and ease of comparison, the evaluation outcomes of DeepCRE alongside traditional methodologies were depicted in an XY plot (Fig. 5c). Notably, the analysis delineated the drug candidates into four distinct sets (A, B, C, and D). Sets A and B comprised drugs favored by DeepCRE predictions but with limited clinical trial evidence, whereas Sets C and D included drugs that were either approved or had substantial clinical trial involvement but were underestimated by DeepCRE predictions.